# DBgDel: Database-Enhanced Gene Deletion Framework for Growth-Coupled Production in Genome-Scale Metabolic Models

### v1

@@ IV-B Results and Performance Comparison
TABLE XII: The performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, and gDel_minRN in the situation of targeting non-extracellular space metabolites of iML1515. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GDLS optGene gDel_minRN
Success rate 507/991 0/991 0/991 508/991
Avg.elapsed time 431.75 244.58 2193.45 2541.23
TABLE XIII: The performance comparison between the proposed method DBgDel with two different initial remaining gene pool settings for iML1515. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GE gene $G_{\text{remain}}$
Success rate 507/991 508/991 508/991
Time.Success = 1 92.46 481.66 90.30
Time.Success = 0 787.17 4561.56 858.01
Time.All cases 431.75 2470.15 463.70
@@ V Discussion and Conclusion
In this study, we developed a novel framework DBgDel to calculate gene deletion strategies to achieve growth-coupled production in genome-scale metabolic models. DBgDel integrates the gene deletion strategy database information to accelerate the algorithms’ computational efficiency. The experimental results above demonstrate the efficiency of the proposed framework, DBgDel, in calculating gene deletion strategies for models of varying scales. Particularly, as the scale of the model increases, its computational efficiency advantages become more pronounced compared to other methods, with an average 6.1-fold acceleration. Notably, even when compared to the previously best-performing method, gDel_minRN, DBgDel achieves a 1.45/2.4/5.89 fold improvement in computational efficiency across three metabolic models, respectively.
@@
Of greater significance, this enhancement in computational efficiency is accompanied by a respectable overall success rate in calculating gene deletion strategies for new target metabolites. Moreover, as the scale of the models increases, the discrepancy in success rates tends to diminish. Specifically, compared to the original gDel_minRN, the gap in success rates for DBgDel across three different models is only 10%/5%/0.1%. The results of the computational experiments affirm that the proposed framework, DBgDel, effectively strikes a balance between success rate and computational efficiency. In further work, we will discuss how to better extract and utilize the information from the gene deletion database, for example, to achieve more flexed information extraction without unified handcrafted computing rules.

### v3

@@ IV-B Results and Performance Comparison
TABLE XII: The performance comparison between the proposed method DBgDel (using Predicted-$G_{\text{remain}}$ genes as the initial remaining gene pool) with GDLS, optGene, gDel_minRN, and gMCSE in the situation of targeting non-extracellular space metabolites of iML1515. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GDLS gMCSE optGene gDel_minRN
Success rate 507/991 0/991 0/991 0/991 508/991
Avg. time 431.75 244.58 1892.39 2193.45 2541.23
TABLE XIII: The performance comparison between the proposed method DBgDel with three different initial remaining gene pool settings for iML1515. Each gene deletion strategy was deemed successful if it achieved a minimum GR and PR of 0.001 or higher during GR maximization. The unit of elapsed time is second (s).
@@
GE gene $G_{\text{remain}}$ RC gene
Success rate 507/991 508/991 508/991 178/991
Time.Success = 1 92.46 481.66 90.30 80.26
Time.Success = 0 787.17 4561.56 858.01 852.43
Time.All cases 431.75 2470.15 463.70 713.74
TABLE XIV: Distribution of percentile ranks for target metabolites PRs across three metabolic models using DBgDel, with MetNetComp PRs as baselines.
Percentile range e_coli_core iMM904 iML1515
0% 8.82% 25.56% 0.91%
0-30% 38.23% 6.67% 1.21%
30-60% 0.00% 4.44% 31.72%
60-90% 0.00% 1.11% 0.91%
>90% 52.94% 62.22% 65.26%
Comparison of PR Percentile Ranks: To further analyze the trade-offs in DBgDel and discuss the limitations of its resulting gene deletion solutions, we compare the PRs obtained by DBgDel with the PRs listed in MetNetComp for target metabolites in three models under study. The DBgDel PRs were derived from computational simulations of metabolic models using DBgDel gene deletion strategies. The MetNetComp PRs were obtained from the MetNetComp database as a baseline for comparison.
Table XIV summarizes the resulting distribution of percentile ranks for target metabolites PRs across three metabolic models using DBgDel. Percentile ranks were calculated using MetNetComp PRs as baselines. A percentile of 0% indicates a zero production rate and, consequently, a failure to achieve growth-coupled production. In the e_coli_core model, the largest proportion of target metabolites (52.94%) fell within the >90% percentile range, followed by 38.23% in the 0-30% range. Smaller proportions were observed in the 0% (8.82%) and the 30-60% and 60-90% ranges (both 0%). For the iMM904 model, the largest proportion of target metabolites (62.22%) were in the >90% range, followed by 25.56% in the 0% range. Smaller proportions were observed in the 0-30% (6.67%), 30-60% (4.44%), and 60-90% (1.11%) ranges. In the iML1515 model, the >90% range had the highest proportion of metabolites (65.26%), followed by 31.72% in the 30-60% range. The 0% range accounted for 0.91%, and the 0-30% and 60-90% ranges accounted for 1.21% and 0.91%, respectively.
@@ V Discussion and Conclusion
In this study, we developed a novel framework DBgDel to calculate gene deletion strategies to achieve growth-coupled production in genome-scale metabolic models. DBgDel integrates the gene deletion strategy database information to accelerate the algorithms’ computational efficiency. The experimental results above demonstrate the efficiency of the proposed framework, DBgDel, in calculating gene deletion strategies for models of varying scales. Particularly, as the scale of the model increases, its computational efficiency advantages become more pronounced compared to other methods, with an average 6.1-fold acceleration. Notably, even when compared to the previously best-performing method, i.e., gDel_minRN, DBgDel achieves a 1.45/2.4/5.89 fold improvement in computational efficiency across three metabolic models, respectively.
@@
Of greater significance, this enhancement in computational efficiency is accompanied by a respectable overall success rate in calculating gene deletion strategies for new target metabolites. Moreover, as the scale of the models increases, the discrepancy in success rates tends to diminish. Specifically, compared to the original gDel_minRN, the gap in success rates for DBgDel across three different models is only 10%/5%/0.1%. The results of the computational experiments affirm that the proposed framework, DBgDel, effectively strikes a balance between success rate and computational efficiency.
We conducted a small-scale experiment using the e_coli_core model to evaluate the integration of a baseline method, gMCSE, into the proposed framework. The results showed a 32% improvement in computational speed while the success rate decreased from 17.5% to 12.5%. The robust trade-off between success rate and computation time was observed. In this study, we configured the default settings of the gMCSE algorithm to ensure a logical setup, requiring that the designed strains produce metabolites during growth and meet the conditions for directional growth coupling. Additionally, we adhered to the framework’s default initial gene pool, the Predicted-$G_{remain}$ gene set. While the application of other algorithms and configurations to large-scale metabolic models remains unexplored, this example demonstrates that the proposed framework, with its initial gene pool, can be adapted to other gene deletion algorithms that properly account for GPR rules.
It is important to note that the implementation of GPR rules in this study differs from that in the gMCSE study described in [33] and [22]. Our method directly implements Boolean functions, whereas the gMCSE method transforms GPR rules into a stoichiometric representation, resulting in some approximation. Specifically, in the gMCSE study, when a reaction is not repressed by genes, its corresponding reaction rate cannot be 0. In contrast, in our study, even if a reaction is not repressed by genes, its reaction rate can still be 0. As a result, gene deletions identified by gMCSE may not always be regarded as successful in our study particularly when a gene is associated with multiple reactions. Additionally, cases where the algorithms reach the computational time limit without producing gene deletion results are also treated as unsuccessful.
Algorithms capable of calculating gene deletion strategies with a maximum number of gene knockouts are desirable. Such algorithms provide gene deletion strategies that reveal the core flow in the metabolic networks. In other words, these algorithms help biological understanding of what is necessary for growth-coupled production, holding significant practical relevance in metabolic engineering. Moreover, based on the resulting maximal gene deletion strategies, it is also possible to derive many other gene deletion strategies with fewer gene deletions [34]. Nowadays, researchers have successfully utilized gene deletion algorithms to address real-world challenges in constructing mutant strains for achieving growth-coupled production of various target metabolites, with applications in biofuel production and industrial production of platform chemicals [35, 22, 23].
The experimental results using the RC genes indicate that STEP 1 of DBgDel effectively narrows the search space. The differences in success rates are 7%, 9%, and 33% for e_coli_core, iMM904, and iML1515, respectively, demonstrating that the effect becomes more pronounced as the model size increases.
The results of the comparison of PR percentile ranks demonstrate that, although DBgDel makes trade-offs between success rate and computational efficiency, its resulting gene deletion strategies still achieve high PRs (particularly in the >90% percentile range) for a significant portion of target metabolites across all three metabolic models when compared to MetNetComp baselines. This also implies that even compared to global optimization methods like RobustKnock [36], —which cannot effectively compute large-scale gene deletions for genome-scale metabolic models—the iterative approach of DBgDel still delivers competitive local solutions in terms of resulting PRs.
The gene deletion strategies of MetNetComp achieve a local maximum in terms of the number of gene deletions. In contrast, the strategies generated by DBgDel are ensured to be neither local nor global maxima, as valid solutions may be obtained by adding or removing genes. Rather than optimizing the number of gene deletions or PR, DBgDel aims to identify acceptable gene deletion strategies efficiently within a short time.
In future work, we will explore reducing the search space for other gene deletion algorithms, demonstrating the broader applicability and potential performance gains of the DBgDel framework. Additionally, we will study how to better extract and utilize information from the gene deletion database, such as achieving more flexible information extraction without relying on unified, handcrafted computing rules.